Consider the function, f(x)=2x³ - 24x² -7. Find the intervals of x where f(x) is


increasing or decreasing.

1
Expert's answer
2022-03-14T12:32:30-0400

Given , f(x) = 2x3 - 24x2 - 7

At the critical point , we know that

f'(x) = 0

So, differentiate the given equation with respect x

f'(x) = 6x2 - 48x = 0

6x(x- 8) = 0

6x = 0 , x-8=0

x=0, x =8

1) f'(x) > 0 , if x<0 and x>8 So, f(x) increases in (-"\infin", 0) U (7, "\infin")

2). f'(x) < 0 , if 0<x<8

So, f(x) decreases in (0,8)